# Random Spiking Neural Networks are Stable and Spectrally Simple

### v1

@@ C.2 Proof of Theorem 2
Let us denote $\overline{x}_{T}=\frac{1}{T}\sum_{k=1}^{T}x_{k}$, $\overline{y}_{T}=\frac{1}{T}\sum_{k=1}^{T}y_{k}$ and $d_{H}(x_{t},y_{t})=\lfloor\nu_{t}n\rfloor$ with $\nu_{t}\in[0,1]$. Let us define $\nu:=\max_{t\in[T]}\ \nu_{t}$ and assume $\nu=\mathcal{O}(\frac{1}{\sqrt{n}})$. We follow a similar strategy to that of Jonasson et al. (2023). Notice that, for all $l\in[L]$, the probability that $s^{(l)}_{T,i}(x_{1},\dots,x_{T},W)$ and $s^{(l)}_{T,i}(y_{1},\dots,y_{T},W)$ differs depends only on the number of neurons that have at least one disagreement at any time at layer l1l-1 and not on where they disagree. We define $D^{(l)}_{T}$ as the number of neurons at layer $l$ that have at least one disagreement at any time, that is $D^{(l)}_{T}:=\frac{1}{2}\sum_{i=1}^{n}\max_{k\in[T]}|s^{(l)}_{k,i}(x_{1},\dots,x_{k},W)-s^{(l)}_{k,i}(y_{1},\dots,y_{k},W)|$. With this, $D^{(l)}_{T}$ is a Markov chain with $n+1$ states, where $D^{(l)}_{T}=0$ is an absorbing state. More precisely, let us denote with $M_{i}:=\frac{1}{2}\max_{k\in[T]}|s^{(l)}_{k,i}(x_{1},\dots,x_{k},W)-s^{(l)}_{k,i}(y_{1},\dots,y_{k},W)|$. Notice that $M_{1},\dots,M_{n}$ are independent random variables with value in $\{0,1\}$. In particular, exploiting Theorem 1, it holds that

### v2

@@ C.2 Proof of Theorem 2
Let us denote $\overline{x}_{T}=\frac{1}{T}\sum_{k=1}^{T}x_{k}$, $\overline{y}_{T}=\frac{1}{T}\sum_{k=1}^{T}y_{k}$ and $d_{H}(x_{t},y_{t})=\lfloor\nu_{t}n\rfloor$ with $\nu_{t}\in[0,1]$. Let us define $\nu:=\max_{t\in[T]}\ \nu_{t}$ and assume $\nu=\mathcal{O}(\frac{1}{\sqrt{n}})$. We follow a similar strategy to that of jonasson2023noise. Notice that, for all $l\in[L]$, the probability that $s^{(l)}_{T,i}(x_{1},\dots,x_{T},W)$ and $s^{(l)}_{T,i}(y_{1},\dots,y_{T},W)$ differs depends only on the number of neurons that have at least one disagreement at any time at layer l1l-1 and not on where they disagree. We define $D^{(l)}_{T}$ as the number of neurons at layer $l$ that have at least one disagreement at any time, that is $D^{(l)}_{T}:=\frac{1}{2}\sum_{i=1}^{n}\max_{k\in[T]}|s^{(l)}_{k,i}(x_{1},\dots,x_{k},W)-s^{(l)}_{k,i}(y_{1},\dots,y_{k},W)|$. With this, $D^{(l)}_{T}$ is a Markov chain with $n+1$ states, where $D^{(l)}_{T}=0$ is an absorbing state. More precisely, let us denote with $M_{i}:=\frac{1}{2}\max_{k\in[T]}|s^{(l)}_{k,i}(x_{1},\dots,x_{k},W)-s^{(l)}_{k,i}(y_{1},\dots,y_{k},W)|$. Notice that $M_{1},\dots,M_{n}$ are independent random variables with value in $\{0,1\}$. In particular, exploiting Theorem 1, it holds that